OpenStudy (anonymous):

HELP PLEASE Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (anonymous):

Use the point-slope form\[y-y_1=m(x-x_1)\]So\[y+4=-\frac{1}{3}(x-6)\]

OpenStudy (anonymous):

If you need to get it to slope intercept form, distribute the slope across the sum on the right side and isolate the y. Like this\[y+4=-\frac{1}{3}(x-6) \implies y=-\frac{x}{3}-2\]

OpenStudy (anonymous):

thank you that helped alot i got it now :)